180 chocolates are shared by Anna and Sara in the ratio 2:3. How much chocolate does each girl receive?
Answer by ikleyn(53765)

.

Anna has 2x chocolates; Sara has 3x chocolate.


The "x" is an unknown number you want to determine.


    2x + 3x = 180

    5x      = 180

     x      = 180/5 = 36.


ANSWER  Anna has  2*36 = 72 chocolates.  Sara has 3*36 = 108 chocolates.

Solved.